What are the coordinates of the point 3 units from the y-axis and at distance square root of 5 from (5,3)?
Answer by lwsshak3(11628)

What are the coordinates of the point 3 units from the y-axis and at distance square root of 5 from (5,3)?
**
x-coordinate 3 units from the y-axis=3
solving for y-coordinate:
distance between (3,y) and (5,3)=√5
using distance formula:
√5=√[(5-3)^2+(3-y)^2]
√5=√[4+9-6y+y^2]
square both sides
5=13-6y+y^2
y^2-6y+8=0
(y-2)(y-4)=0
y=2
or
y=4
Coordinates are:
(3,2) or (3,4)